OpenStudy (anonymous):

6(5.3 X 10^-4) Can someone do this and show their work for me to help me understand?

OpenStudy (anonymous):

\[6(5.3 x 10^{-4)}\]

OpenStudy (anonymous):

Use PEMDAS P= Parenthesis E=Exponents M=Multiply D= Divide A= Add S= Subtract

OpenStudy (anonymous):

Pemdas doesn't help me, trust me i spent three hours yesterday trying to figure this out

OpenStudy (anonymous):

HELP ME.

ganeshie8 (ganeshie8):

hey wat do you wanto to do here ?

ganeshie8 (ganeshie8):

\(\large 6(5.3 \times 10^{-4}) \)

ganeshie8 (ganeshie8):

if you wanto simplify, u just need to multiply 6 and 5.3 together

ganeshie8 (ganeshie8):

\(\large 6(5.3 \times 10^{-4}) \) \(\large 31.8\times 10^{-4} \)

ganeshie8 (ganeshie8):

you wanto put it in scientific notation ?

ganeshie8 (ganeshie8):

then, it becomes :- \(\large 6(5.3 \times 10^{-4}) \) \(\large 31.8\times 10^{-4} \) \(\large 3.18\times 10^{-3} \)

OpenStudy (anonymous):

no i know the answer, its 3.18 X 10^-3

ganeshie8 (ganeshie8):

nice :) which part is throwing u off ? :)

OpenStudy (anonymous):

nevermind, think i got it now, thanks (:

ganeshie8 (ganeshie8):

remember this :- when u move the decimal left by 1, add 1 to exponent

ganeshie8 (ganeshie8):

since we moved the decimal to left by 1 unit, we have added 1 to exponent -4 -4+1 = -3

OpenStudy (anonymous):

Could you help with one more?

ganeshie8 (ganeshie8):

sure

OpenStudy (anonymous):

Find the balance in the given account after the given period. $500 principle earning 4% compounded quarterly, after 6 years.

ganeshie8 (ganeshie8):

use compound interest formula :- \(\large A = P\left(1+\frac{r}{n}\right)^{n\times t}\) \(P\) = Princial to start wid(500) \(r\) = rate of interest(4% = 0.04) \(n\) = compounding interval(quarterly = 4) \(t\) = time in years (6)

ganeshie8 (ganeshie8):

plug and chug

OpenStudy (anonymous):

Okay

ganeshie8 (ganeshie8):

wat do you get ? :)

OpenStudy (anonymous):

No lol

ganeshie8 (ganeshie8):

\(\large A = 500\left(1+\frac{0.04}{4}\right)^{4\times 6} \)

ganeshie8 (ganeshie8):

use ur calculator

OpenStudy (anonymous):

I'm lost

ganeshie8 (ganeshie8):

you need to use ur calculaotor, and find that value

ganeshie8 (ganeshie8):

wolfram says balance after 6 years wud be 634.9 dollars http://www.wolframalpha.com/input/?i=500%281%2B.04%2F4%29%5E%284*6%29

OpenStudy (anonymous):

whats 1.01 to the 25th power?

ganeshie8 (ganeshie8):

is that a different question ?

OpenStudy (anonymous):

noo.. its part of this one, i just dont know how to do that in a calculator

ganeshie8 (ganeshie8):

ohk... got u :) its actually 1.01 to the 24th power right ?

OpenStudy (anonymous):

Opps yes

ganeshie8 (ganeshie8):

cuz, in the exponent 4x6 = 24 its NOT 25

ganeshie8 (ganeshie8):

:) ok, i can teach u how to plug exponents in ur calculator which calculator u using ?

OpenStudy (anonymous):

i know lol its 24

OpenStudy (anonymous):

idk,? lol

ganeshie8 (ganeshie8):

i mean, which calculator u using...

ganeshie8 (ganeshie8):

do u have a calculator ?

OpenStudy (anonymous):

a scientific one?

ganeshie8 (ganeshie8):

it must be having a name

ganeshie8 (ganeshie8):

do u have windows calculator, in ur computer ?

OpenStudy (anonymous):

noooo

ganeshie8 (ganeshie8):

ohk, here is an online calculator site :- http://www.wolframalpha.com

ganeshie8 (ganeshie8):

go there, and enter :- ``` 1.01^24 ```

ganeshie8 (ganeshie8):

wat do u get ?

OpenStudy (anonymous):

1.27?

ganeshie8 (ganeshie8):

good :)

OpenStudy (anonymous):

so 1.27x500?

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

Okay(:

ganeshie8 (ganeshie8):

good job !!

OpenStudy (anonymous):

thank you(:

ganeshie8 (ganeshie8):

np :)